OpenStudy (anonymous):

8x + 5y = 3 7x – 3y = 10 can you please answer in elimination form?

OpenStudy (anonymous):

What have you tried so far? Show all of your work.

OpenStudy (anonymous):

Use Elimination Method..)

OpenStudy (anonymous):

7 (8x + 5y = 3) 8 (7x – 3y = 10 )

OpenStudy (anonymous):

= 56x + 35 y = 21 56x - 24y = 80 hope...this helps

OpenStudy (anonymous):

nw Subtract Both eq and find y then equate it to any eq and find x